# Quantum Zeno Effect in Noisy Integrable Quantum Circuits for Impurity Models

### v1

@@ V N-particle Bethe Ansatz
$H=\sum_{j=1}^{L-1}\sum_{\sigma=\uparrow,\downarrow}(c^{\dagger}_{j,\sigma}c_{j+1,\sigma}+c^{\dagger}_{j+1,\sigma}c_{j,\sigma})+J(c^{\dagger}_{0,\sigma}c_{1,\sigma}+c^{\dagger}_{1,\sigma}c_{0,\sigma})+4i\gamma\sum_{j=1}^{L}n_{j,\uparrow}n_{j,\downarrow}-2i\gamma\sum_{j=1}^{L}\sum_{\sigma=\uparrow,\downarrow}n_{j,\sigma}.$ (S.60)
@@
$S_{b_{j}}(k)=\frac{-e^{2ik_{j}}J^{2}+e^{2ik_{j}}+1}{-J^{2}+\left(1+e^{2ik_{j}}\right)}=e^{2ik_{J}}\frac{2\cos(k_{j})-iJ^{2}\sin(k_{j})-J^{2}\cos(k_{j})}{2\cos(k_{j})+iJ^{2}\sin(k_{j})-J^{2}\cos(k_{j})},$ (S.61)
@@
$S_{j,l}=\frac{\sin(k_{j})-\sin(k_{l})-2\gamma P_{j,l}}{\sin(k_{j})-\sin(k_{l})-2\gamma},$ (S.62)
@@
Now, the usual Bethe Ansatz routine can be employed. It is very important to notice that the boundary $S-$matrix does not depend on the spin variable but the bulk $S-$matrix depends on the spin variable (through the permutation operator). The Yang-Baxter equation is therefore trivially satisfied, thereby proving that the model is integrable.
@@
Then for the open boundary condition, we consider the double-row transfer matrix
@@
$\Xi(\lambda)=T_{0}(\lambda)\hat{T}(\lambda).$ (S.64)
@@
$t(\lambda)=\operatorname{tr}_{0}\Xi(\lambda).$ (S.65)
@@
$[t(\lambda),t(\rho)]=0.$ (S.66)
First notice that using coordinate Bethe Ansatz, one can immediately construct the eigenvalue problem for the transfer matrix
$\tau_{j}=$ $S_{j-1,j}\left(k_{j-1},k_{j}\right)\cdots S_{1,j}\left(k_{1},k_{j}\right){S_{b_{j}}}\left(k_{j}\right)S_{j,1}\left(-k_{j},k_{1}\right)S_{j,j-1}\left(-k_{j},k_{j-1}\right)$ (S.67)
@@
The eigenvalue of the transfer matrix satisfies the $T-Q$ relation
@@
$\quad+\prod_{j=1}^{N}\frac{2\lambda}{2\lambda+\eta}\left(\lambda-\sin k_{j}\right)\left(\lambda+\sin k_{j}\right)\frac{Q(\lambda+\eta)}{Q(\lambda)},$ (S.69)
@@
$Q(\lambda)=\prod_{l=1}^{M}(\lambda-\lambda_{l})(\lambda+\lambda_{l}+\eta).$ (S.70)
Imposing the regularity condition of the $T-Q$ relation immediately gives the Bethe equation
$\frac{\lambda_{j}+\eta}{\lambda_{j}}\prod_{j=1}^{N}\frac{\lambda_{j}-\sin(k_{j})+\eta}{\lambda_{j}-\sin(k_{j})}\frac{\lambda_{j}+\sin(k_{j})+\eta}{\lambda_{j}+\sin(k_{j})}=-\prod_{l=1}^{M}\frac{\lambda_{j}-\lambda_{l}+\eta}{\lambda_{j}-\lambda_{l}-\eta}\frac{\lambda_{j}+\lambda_{l}+2\eta}{\lambda_{j}+\lambda_{l}}.$ (S.71)
@@
$\frac{\lambda_{j}+\frac{\eta}{2}}{\lambda_{j}-\frac{\eta}{2}}\prod_{j=1}^{N}\frac{\lambda_{j}-\sin(k_{j})+\frac{\eta}{2}}{\lambda_{j}-\sin(k_{j})-\frac{\eta}{2}}\frac{\lambda_{j}+\sin(k_{j})+\frac{\eta}{2}}{\lambda_{j}+\sin(k_{j})-\frac{\eta}{2}}=-\prod_{l=1}^{M}\frac{\lambda_{j}-\lambda_{l}+\eta}{\lambda_{j}-\lambda_{l}-\eta}\frac{\lambda_{j}+\lambda_{l}+\eta}{\lambda_{j}+\lambda_{l}-\eta}.$ (S.72)
From Eq. (S.68), upon using the shift $\lambda\to\lambda-\eta/2$ we obtain
$e^{-2ik_{j}L}/S_{b_{j}}(k_{j})=\prod_{l=1}^{M}\frac{\sin k_{j}-\lambda_{l}+\frac{\eta}{2}}{\sin k_{j}-\lambda_{l}-\frac{\eta}{2}}\frac{\sin k_{j}+\lambda_{l}+\frac{\eta}{2}}{\sin k_{j}+\lambda_{l}-\frac{\eta}{2}}.$ (S.73)
@@
$e^{-2ik_{J}(L+1)}\frac{2\cos(k_{j})+iJ^{2}\sin(k_{j})-J^{2}\cos(k_{j})}{2\cos(k_{j})-iJ^{2}\sin(k_{j})-J^{2}\cos(k_{j})}=\prod_{l=1}^{M}\frac{\sin k_{j}-\lambda_{l}+\frac{\eta}{2}}{\sin k_{j}-\lambda_{l}-\frac{\eta}{2}}\frac{\sin k_{j}+\lambda_{l}+\frac{\eta}{2}}{\sin k_{j}+\lambda_{l}-\frac{\eta}{2}}.$ (S.74)
@@
$e^{2ik_{j}(L+1)}\frac{2\cos(k_{j})-iJ^{2}\sin(k_{j})-J^{2}\cos(k_{j})}{2\cos(k_{j})+iJ^{2}\sin(k_{j})-J^{2}\cos(k_{j})}=\prod_{l=1}^{M}\frac{\sin k_{j}-\lambda_{l}+\gamma}{\sin k_{j}-\lambda_{l}-\gamma}\frac{\sin k_{j}+\lambda_{l}+\gamma}{\sin k_{j}+\lambda_{l}-\gamma},$ (S.75)
@@
$\frac{\lambda_{j}+\gamma}{\lambda_{j}-\gamma}\prod_{j=1}^{N}\frac{\lambda_{j}-\sin(k_{j})+\gamma}{\lambda_{j}-\sin(k_{j})-\gamma}\frac{\lambda_{j}+\sin(k_{j})+\gamma}{\lambda_{j}+\sin(k_{j})-\gamma}=-\prod_{l=1}^{M}\frac{\lambda_{j}-\lambda_{l}+2\gamma}{\lambda_{j}-\lambda_{l}-2\gamma}\frac{\lambda_{j}+\lambda_{l}+2\gamma}{\lambda_{j}+\lambda_{l}-2\gamma}.$ (S.76)
@@
If one is interested in the many-point correlation function, one needs to solve the many-particle BAE given by Eq. (S.75) and Eq. (S.76). In the main text, we only considered the two-point correlation function, and we leave the general problem of the 2n-point correlation function for future work. The Bethe equations for the real Hubbard interaction were first obtained in  [84].
@@
Figure S.4: The magnetization $S^{z}_{j}(t)$ for time $t$ and site $j$ in the XXX case with system size $L=10$ (right panel) and XX case with system size $L=40$ (left panel) for dephasing rate $\gamma=3$.
VI Numerical details
Here we will show the detailed calculation to get the diffusion constant $D$ numerically for the Heisenberg (XXX) and XX case. We start from a Doman wall configuration initially with $\rho_{D}=\sum_{z\in\mathcal{S}_{D}}\sigma^{+}_{z}\ket{0}\bra{0}\sigma^{+}_{z}$ and $\mathcal{S}_{D}=0,1,2,...,L/2$. Then we evolve the state $\rho(t)$ via the Lindbladian master equation (S.19) from the initial state $\rho(0)=\rho_{D}$. Then we calculated the magnetization profile $S^{z}_{j}(t)$ for each time $t$ and site $j$. The results for $\gamma=3$ in the XX case and XXX case are shown in Fig. S.4. For the XXX case, the domain wall hit the boundary when $t\sim 50$ and resulting in a finite size effect.
@@ VI Numerical details
If the bulk is diffusive, then in the limit lattice spacing $a\to 0$ and system size $L\to\infty$ $S^{z}_{j}(t)\to S^{z}(x,t)$ should satisfy the diffusion equation $(\partial_{t}-D\partial_{x}^{2})S^{z}(x,t)=0$. The initial condition is a domain wall meaning $S^{z}(x,0)=\mathrm{Sign}(L/2-x)$ and the solution reads $S^{z}(x,t)=-\mathrm{Erf}(\frac{x}{\sqrt{Dt}})$. This is obvious that for different times and spaces, the function collapses on the curve with a normalized variable $\xi=x/\sqrt{t}$ in the way that $S^{z}(\xi)=-\mathrm{Erf}(\frac{\xi}{\sqrt{D}})$. The system is not perfectly diffusive for early-time dynamics and due to the finite size effect, we cannot evolve the system for an extremely long time. Therefore, we will fit the magnetization $S^{z}(\xi)$ to $-\mathrm{Erf}(\xi/\sqrt{D(t)})$ for a finite time $t$ to get the diffusion constant $D(t)$. To get the long-time behavior, we expand $D(t)=D(\infty)+\frac{D_{1}}{t}+\mathcal{O}(\frac{1}{t^{2}})$ and scale the diffusion constant when $t\to\infty$. The fitting and scaling are shown in Fig.S.5. We found for both XX and XXX case $D_{0}$ which is the diffusion constant for long-time dynamics is $D_{0}=\frac{1}{2\gamma}$ for big gamma. As for small $\gamma$, due to the finite size effect, we are not able to conclude for the XXX case. For the XX case, where we can go for larger system sizes, we can show $D_{0}=\frac{1}{2\gamma}$.
@@
Figure S.5: Scaling of the diffusion constant $D$ with time $t$ and $\gamma$ for the XXX circuit. The left panel shows $D$ as a fit parameter obtained by fitting the spin profile to the $S^{z}(\xi)=\mathrm{Erf}(\xi/\sqrt{D(t)})$. The right panel shows $D(t)$ as a function for different times and we scale $D(\infty)$ by taking $1/t\to 0$. The dashed line is the fitting for $1/t\to 0$ and different colors for different $\gamma$.

### v2

@@ Appendix E N-particle Bethe Ansatz
$H=\sum_{j=1}^{L-1}\sum_{\sigma=\uparrow,\downarrow}(c^{\dagger}_{j,\sigma}c_{j+1,\sigma}+c^{\dagger}_{j+1,\sigma}c_{j,\sigma})+J(c^{\dagger}_{0,\sigma}c_{1,\sigma}+c^{\dagger}_{1,\sigma}c_{0,\sigma})+4i\gamma\sum_{j=1}^{L}n_{j,\uparrow}n_{j,\downarrow}-2i\gamma\sum_{j=1}^{L}\sum_{\sigma=\uparrow,\downarrow}n_{j,\sigma}.$ (69)
@@
$S_{b_{j}}(k)=\frac{-e^{2ik_{j}}J^{2}+e^{2ik_{j}}+1}{-J^{2}+\left(1+e^{2ik_{j}}\right)}=e^{2ik_{J}}\frac{2\cos(k_{j})-iJ^{2}\sin(k_{j})-J^{2}\cos(k_{j})}{2\cos(k_{j})+iJ^{2}\sin(k_{j})-J^{2}\cos(k_{j})},$ (70)
@@
$S_{j,l}=\frac{\sin(k_{j})-\sin(k_{l})-2\gamma P_{j,l}}{\sin(k_{j})-\sin(k_{l})-2\gamma},$ (71)
@@
Now, the usual Bethe Ansatz routine can be employed. It is very important to notice that the boundary $S$-matrix does not depend on the spin variable, but the bulk $S$-matrix depends on the spin variable (through the permutation operator). The Yang-Baxter equation is, therefore, trivially satisfied, thereby proving that the model is integrable.
@@
Then, for the open boundary condition, we consider the double-row transfer matrix
@@
$\Xi(\lambda)=T_{0}(\lambda)\hat{T}_{0}(\lambda).$ (73)
@@
$t(\lambda)=\operatorname{tr}_{0}\Xi(\lambda).$ (74)
@@
$t(\lambda),t(\rho)]=0.$ (75)
First, notice that using the coordinate Bethe Ansatz, one can immediately construct the eigenvalue problem for the transfer matrix
$\tau_{j}=$ $S_{j-1,j}\left(k_{j-1},k_{j}\right)\cdots S_{1,j}\left(k_{1},k_{j}\right){S_{b_{j}}}\left(k_{j}\right)S_{j,1}\left(-k_{j},k_{1}\right)S_{j,j-1}\left(-k_{j},k_{j-1}\right)$ (76)
@@
The eigenvalue of the transfer matrix satisfies the TQT-Q relation
@@
$\quad+\prod_{j=1}^{N}\frac{2\lambda}{2\lambda+\eta}\left(\lambda-\sin k_{j}\right)\left(\lambda+\sin k_{j}\right)\frac{Q(\lambda+\eta)}{Q(\lambda)},$ (78)
@@
$Q(\lambda)=\prod_{l=1}^{M}(\lambda-\lambda_{l})(\lambda+\lambda_{l}+\eta).$ (79)
Imposing the regularity condition of the TQT-Q relation immediately gives the Bethe equation
$\frac{\lambda_{j}+\eta}{\lambda_{j}}\prod_{j=1}^{N}\frac{\lambda_{j}-\sin(k_{j})+\eta}{\lambda_{j}-\sin(k_{j})}\frac{\lambda_{j}+\sin(k_{j})+\eta}{\lambda_{j}+\sin(k_{j})}=-\prod_{l=1}^{M}\frac{\lambda_{j}-\lambda_{l}+\eta}{\lambda_{j}-\lambda_{l}-\eta}\frac{\lambda_{j}+\lambda_{l}+2\eta}{\lambda_{j}+\lambda_{l}}.$ (80)
@@
$\frac{\lambda_{j}+\frac{\eta}{2}}{\lambda_{j}-\frac{\eta}{2}}\prod_{j=1}^{N}\frac{\lambda_{j}-\sin(k_{j})+\frac{\eta}{2}}{\lambda_{j}-\sin(k_{j})-\frac{\eta}{2}}\frac{\lambda_{j}+\sin(k_{j})+\frac{\eta}{2}}{\lambda_{j}+\sin(k_{j})-\frac{\eta}{2}}=-\prod_{l=1}^{M}\frac{\lambda_{j}-\lambda_{l}+\eta}{\lambda_{j}-\lambda_{l}-\eta}\frac{\lambda_{j}+\lambda_{l}+\eta}{\lambda_{j}+\lambda_{l}-\eta}.$ (81)
From Eq. (77), upon using the shift $\lambda\to\lambda-\eta/2$ we obtain
$e^{-2ik_{j}L}/S_{b_{j}}(k_{j})=\prod_{l=1}^{M}\frac{\sin k_{j}-\lambda_{l}+\frac{\eta}{2}}{\sin k_{j}-\lambda_{l}-\frac{\eta}{2}}\frac{\sin k_{j}+\lambda_{l}+\frac{\eta}{2}}{\sin k_{j}+\lambda_{l}-\frac{\eta}{2}}.$ (82)
@@
$e^{-2ik_{J}(L+1)}\frac{2\cos(k_{j})+iJ^{2}\sin(k_{j})-J^{2}\cos(k_{j})}{2\cos(k_{j})-iJ^{2}\sin(k_{j})-J^{2}\cos(k_{j})}=\prod_{l=1}^{M}\frac{\sin k_{j}-\lambda_{l}+\frac{\eta}{2}}{\sin k_{j}-\lambda_{l}-\frac{\eta}{2}}\frac{\sin k_{j}+\lambda_{l}+\frac{\eta}{2}}{\sin k_{j}+\lambda_{l}-\frac{\eta}{2}}.$ (83)
@@
$e^{2ik_{j}(L+1)}\frac{2\cos(k_{j})-iJ^{2}\sin(k_{j})-J^{2}\cos(k_{j})}{2\cos(k_{j})+iJ^{2}\sin(k_{j})-J^{2}\cos(k_{j})}=\prod_{l=1}^{M}\frac{\sin k_{j}-\lambda_{l}+\gamma}{\sin k_{j}-\lambda_{l}-\gamma}\frac{\sin k_{j}+\lambda_{l}+\gamma}{\sin k_{j}+\lambda_{l}-\gamma},$ (84)
@@
$\frac{\lambda_{j}+\gamma}{\lambda_{j}-\gamma}\prod_{j=1}^{N}\frac{\lambda_{j}-\sin(k_{j})+\gamma}{\lambda_{j}-\sin(k_{j})-\gamma}\frac{\lambda_{j}+\sin(k_{j})+\gamma}{\lambda_{j}+\sin(k_{j})-\gamma}=-\prod_{l=1}^{M}\frac{\lambda_{j}-\lambda_{l}+2\gamma}{\lambda_{j}-\lambda_{l}-2\gamma}\frac{\lambda_{j}+\lambda_{l}+2\gamma}{\lambda_{j}+\lambda_{l}-2\gamma}.$ (85)
@@
If one is interested in the many-point correlation function, one needs to solve the many-particle BAE given by Eq. (84) and Eq. (85). In the main text, we only considered the two-point correlation function, and we leave the general problem of the 2n-point correlation function for future work. The Bethe equations for the real Hubbard interaction were first obtained in  [91].
@@
Figure 7: The magnetization $S^{z}_{j}(t)$ for time $t$ and site $j$ in the XXX case with system size $L=10$ (right panel) and XX case with system size $L=40$ (left panel) for dephasing rate $\gamma=3$.
Appendix F Numerical details
[…]
@@ Appendix F Numerical details
If the bulk is diffusive, then in the limit lattice spacing $a\to 0$ and system size $L\to\infty$ $S^{z}_{j}(t)\to S^{z}(x,t)$ should satisfy the diffusion equation $(\partial_{t}-D\partial_{x}^{2})S^{z}(x,t)=0$. The initial condition is a domain wall meaning $S^{z}(x,0)=\mathrm{Sign}(L/2-x)$ and the solution reads $S^{z}(x,t)=-\mathrm{Erf}(\frac{x}{\sqrt{Dt}})$. It is obvious that for different times and spaces, the function collapses on the curve with a normalized variable $\xi=x/\sqrt{t}$ in the way that $S^{z}(\xi)=-\mathrm{Erf}(\frac{\xi}{\sqrt{D}})$. The system is not perfectly diffusive for early-time dynamics, and due to the finite-size effect, we cannot evolve the system for an extremely long time. Therefore, we will fit the magnetization $S^{z}(\xi)$ to $\mathrm{Erf}(\xi/\sqrt{D(t)})$ for a finite time $t$ to get the diffusion constant $D(t)$. To get the long-time behavior, we expand $D(t)=D(\infty)+\frac{D_{1}}{t}+\mathcal{O}(\frac{1}{t^{2}})$ and scale the diffusion constant when $t\to\infty$. The fitting and scaling are shown in Fig.8. We found for both the XX and XXX cases, $D_{0}$, which is the diffusion constant for long-time dynamics, is $D_{0}=\frac{1}{2\gamma}$ for big gamma. As for small $\gamma$, due to the finite size effect, we are not able to conclude for the XXX case. For the XX case, where we can go for larger system sizes, we can show $D_{0}=\frac{1}{2\gamma}$.
@@
Figure 8: Scaling of the diffusion constant $D$ with time $t$ and $\gamma$ for the XXX circuit. The left panel shows $D$ as a fit parameter obtained by fitting the spin profile to the $S^{z}(\xi)=\mathrm{Erf}(\xi/\sqrt{D(t)})$. The right panel shows $D(t)$ as a function for different times, and we scale $D(\infty)$ by taking $1/t\to 0$. The dashed line is the fitting for $1/t\to 0$ and different colors for different $\gamma$.